### v3

# VisionReward: Fine-Grained Multi-Dimensional Human Preference Learning for Image and Video Generation

Abstract
Visual generative models have achieved remarkable progress in synthesizing photorealistic images and videos, yet aligning their outputs with human preferences across critical dimensions remains a persistent challenge. Though reinforcement learning from human feedback offers promise for preference alignment, existing reward models for visual generation face limitations, including black-box scoring without interpretability and potentially resultant unexpected biases. We present VisionReward, a general framework for learning human visual preferences in both image and video generation. Specifically, we employ a hierarchical visual assessment framework to capture fine-grained human preferences, and leverages linear weighting to enable interpretable preference learning. Furthermore, we propose a multi-dimensional consistent strategy when using VisionReward as a reward model during preference optimization for visual generation. Experiments show that VisionReward can significantly outperform existing image and video reward models on both machine metrics and human evaluation. Notably, VisionReward surpasses VideoScore by 17.2% in preference prediction accuracy, and text-to-video models with VisionReward achieve a 31.6% higher pairwise win rate compared to the same models using VideoScore. All code and datasets are provided at https://github.com/THUDM/VisionReward.
1 Introduction
Figure 1: Illustration of how VisionReward works for evaluation and optimization of visual generation. a) Evaluation: VisionReward performs comprehensive evaluation through dimension-specific binary visual QA testing, producing human-aligned, fine-grained assessment scores. b) Optimization: VisionReward enable better preference optimization, enhancing multiple key aspects.
Visual generative models, including text-to-image (ding2021cogview; ramesh2021zero; saharia2022photorealistic; rombach2022high; betker2023improving; podell2023sdxl) and text-to-video (hong2022cogvideo; ho2022imagen; villegas2022phenaki; opensora; chen2024videocrafter2; yang2024cogvideox) generation, have recently experienced rapid developments. Through large-scale pretraining, these models can effectively translate textual descriptions into photorealistic images or temporally coherent videos. To further align them with human preferences, reinforcement learning from human feedback (RLHF) (ouyang2022training)—initially introduced in large language models—has recently been adapted to visual generation tasks (xu2023imagereward; he2024videoscore).
A key bottleneck in applying RLHF to visual generation lies in developing effective visual reward models. Recent studies (xu2023imagereward; kirstain2023pick; wu2023human) have explored training reward models to predict human visual preferences, enabling automatic evaluation and preference optimization for visual generative models. For evaluation, reward models function as automated metrics that quantitatively measure the alignment between generated outputs and human preference criteria (li2023agiqa). For optimization, reward models identify reliable directions for improving visual generation models. Essentially, they can provide feedback in reinforcement learning or generate preference pairs, thus reducing dependence on human annotation (black2023training; fan2023dpok; clark2023directly).
Despite recent progress in reward models (RMs) for visual generation, two primary challenges remain: First, lack of interpretability and risk of unexpected bias. Current RMs for visual generation often suffer from limited interpretability. These models inherently involve complex trade-offs among multiple factors, yet their scoring mechanisms lack transparency regarding how such trade-offs are performed. This opacity raises concerns about potential unexpected biases. Though multimodal LLMs like Gemini (team2024gemini) and GPT-4o (achiam2023gpt) enhance interpretability through explainable rating rationales, their general-purpose architectures usually underperform specialized black-box models in fine-grained assessments (chen2024mj). This raises a key dilemma: how to design preference prediction method to be interpretable while maintaining accuracy.
Second, lack of effective reward models for video generation. The rapid development of text-to-video generative models has intensified the demand for video reward models. Although image reward models can assess individual frame quality, their frame-level nature inherently neglects essential temporal dependencies in video sequences. While VideoScore (he2024videoscore) has pioneered direct video evaluation through learnable metrics, it still suffers from limitations such as insufficient accuracy in preference prediction and optimization in video generation.
Contribution. To address these challenges, we propose a general framework VisionReward to build accurate reward models for both image and video generation. VisionReward is trained with two steps: fine-grained visual assessment and interpretable preference learning. First, to capture human visual preferences, we identify nine major dimensions and decouple preferences into 64 fine-grained questions. Second, to ensure interpretable preference learning, we propose to use the classical linear weighting mechanism on the question outcomes. It enables intuitive visualization of each question’s impact.
To apply it as a reward model for visual generation models, we propose a multi-dimensional consistent strategy during preference optimization. The goal of this strategy is to mitigate unintended and unquantifiable biases. Specifically, a pair of visual samples is used for preference optimization (e.g., DPO (wallace2024diffusion)) only if one sample is consistently preferred over the other across all dimensions.
To summarize, we present VisionReward as a general framework for visual preference learning. Empirically, we show that VisionReward makes the following contributions:
  • We design fine-grained multi-dimensional preference annotation and build the most fine-grained dataset which contain 81K samples and 5M binary annotation, which enable the training of VisionReward.
  • For visual preference prediction, VisionReward achieves state-of-the-art performance across multiple benchmarks, while maintaining interpretability via hierarchical diagnostic QA and explicit linear weighting. For instance, VisionReward outperforms VideoScore (he2024videoscore) by 17.2% in accuracy on preference prediction.
  • For visual generation, VisionReward can serve as an effective reward model for preference optimization (e.g., DPO), significantly enhancing the text-to-image and text-to-video models. For example, video generation models with VisionReward achieve a 31.6% higher pairwise win rate compared to the same models using VideoScore.
Dataset Image Video #Samples #Dimensions #Fine-Grained #Annotation
ImageReward (xu2023imagereward) 9K - - 0.1M
Pick-a-Pic (kirstain2023pick) 38K - - 0.6M
HPDv2 (wu2023human) 430K - - 0.8M
RichHF-18K (liang2024rich) 18K 4 - 0.1M
MPS (zhang2024learning) 608K 4 - 2.4M
VideoScore (he2024videoscore) 38K 5 - 0.2M
VisionReward (Ours) 81K 18–20 61–64 5.0M
Table 1: Comparison of dataset of VisionReward and other datasets.
Figure 2: Illustration of fine-grained multi-dimensional design. (Left) For image: 5 dimensions, 18 sub-dimensions, and 61 binary questions. (Right) For video: 9 dimensions, 20 sub-dimensions, and 64 binary questions.
Figure 3: Overall framework of VisionReward. 1) Fine-grained Visual Assessment: fine-tune multimodal LLM to perform binary visual question-answering through hierarchical dimensions. 2) Interpretable Preference Learning: utilize visual QA outputs to predict preferences through linear weighted summation. 3) Multi-Dimensional Preference Optimization: optimization strategy across multiple dimensions.
2 Related Work
Reinforcement Learning from Human Feedback (RLHF) (stiennon2020learning; nakano2021webgpt; ouyang2022training) refers to optimizing models with reinforcement learning based on human feedback, which is also explored in image and video generation.
Preference Learning for Visual Generation. There are many works learning from human preferences, which collect human annotation for text-to-image (xu2023imagereward; kirstain2023pick; wu2023human) and text-to-video (he2024videoscore). Note that existing approaches (zhang2024learning; liang2024rich) have attempted to augment human annotations or expand dimensions of human preferences in visual generation. Different from them, VisionReward defines fine-grained multi-dimensional human preferences with the goal of disentangling distinct factors to decouple human preferences, to build more accurate and interpretable RM.
RLHF for Visual Generation. For visual generation tasks, several works have explored RLHF, optimizing from the gradient (xu2023imagereward; wu2024deep) or using a policy-based RL approach (black2023training; fan2023dpok; clark2023directly). All these methods require a reward model (RM) to provide feedback for online learning. Diffusion-DPO (wallace2024diffusion) has proposed to optimize the diffusion model directly using human-labeled preference data. However, most RLHF methods face the issue of biased-optimization. By employing a multi-dimensional method, VisionReward achieves robust RLHF.
Preliminary for Diffusion-DPO. Given a data distribution $q(x_{0})$, Diffusion models (sohl2015deep; ho2020denoising; song2020score) contains forward process and reverse process. Forward process $q(x_{1:T}|x_{0})$ gradually add noise to the data $x_{0}$ and reverse process $p_{\theta}(x_{0:T})$ learns transitions to recover data. Training diffusion model can be performed by evidence lower bound (kingma2021variational; song2021maximum):
$L_{\mathrm{DM}}=\mathbb{E}_{\boldsymbol{x}_{0},\boldsymbol{\epsilon}\sim\mathcal{N}(0,\boldsymbol{I}),t}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{x}_{t},t\right)\right\|_{2}^{2}\right],$ (1)
with $t\sim\mathcal{U}(0,T)$ and $\boldsymbol{x}_{t}\sim q\left(\boldsymbol{x}_{t}\mid\boldsymbol{x}_{0}\right)$.
Diffusion-DPO (wallace2024diffusion) introduces direct preference optimization based on preference pairs. We denote the “win” and “lose” samples as $x_{0}^{w},x_{0}^{l}$, and the objective is as follows:
$\mathcal{L}(\theta)=-\mathbb{E}_{\left(\boldsymbol{x}_{0}^{w},\boldsymbol{x}_{0}^{l}\right)\sim\mathcal{D},t\sim\mathcal{U}(0,T),\boldsymbol{x}_{t}^{w}\sim q\left(\boldsymbol{x}_{t}^{w}\mid\boldsymbol{x}_{0}^{w}\right),\boldsymbol{x}_{t}^{l}\sim q\left(\boldsymbol{x}_{t}^{l}\mid\boldsymbol{x}_{0}^{l}\right)}$
$\log\sigma\left(-\beta T\omega\left(\lambda_{t}\right)(\right$.
$\left\|\boldsymbol{\epsilon}^{w}-\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{x}_{t}^{w},t\right)\right\|_{2}^{2}-\left\|\boldsymbol{\epsilon}^{w}-\boldsymbol{\epsilon}_{\mathrm{ref}}\left(\boldsymbol{x}_{t}^{w},t\right)\right\|_{2}^{2}$
$\left.\left.-\left(\left\|\boldsymbol{\epsilon}^{l}-\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{x}_{t}^{l},t\right)\right\|_{2}^{2}-\left\|\boldsymbol{\epsilon}^{l}-\boldsymbol{\epsilon}_{\mathrm{ref}}\left(\boldsymbol{x}_{t}^{l},t\right)\right\|_{2}^{2}\right)\right)\right).$ (2)
DPO is ordinarily based on overall preference, which may be biased. VisionReward enables Multi-Dimensional Preference Optimization to enhance it.
3 VisionReward
3.1 VisionReward Annotation
Fine-Grained Design. When evaluating an image or video, human preferences are often a result of the interplay of multiple factors (palmer2013visual; ibarra2017image), necessitating a balance among various considerations. To deconstruct human preferences systematically, we develop a fine-grained multi-dimensional framework, as shown in  Table 1 and  Fig. 2. For each sub-dimension, we set options that vary gradually in degree, and decompose these options into a series of binary questions (Cf.  Tables 16, 17, 18 and 19 in Appendix Section 10 for details).
Dataset Preparation. The dataset for annotation is collected from diverse sources. For images, we sample images from multiple popular datasets, including ImageRewardDB (xu2023imagereward), HPDv2 (wu2023human), and Pick-a-Pic (kirstain2023pick), and obtain 48k images after filtering. For videos, we sample prompts from VidProM (wang2024vidprom) and use multiple text-to-video models to generate videos to annotate. To ensure diversity of prompts, we use Rouge-L (lin-2004-rouge) for initial filtering, follow UniFL (zhang2024uniflimprovestablediffusion) to perform a semantic-based filtering, and use ChatGPT (achiam2023gpt) for data cleaning, finally get 10k prompts. Then we use CogVideoX (yang2024cogvideox), VideoCrafter2 (chen2024videocrafter2) and OpenSora (opensora) to generate 30k videos, sample from Panda-70M (chen2024panda) to get 3k real videos, leading to 33k videos for annotation. More details are provided in Appendix Section 7.1.
Annotation Management. To avoid bias of annotators, our annotation management includes professional management and standard document. Cooperating with a specialized company, we strictly conduct annotation training for annotators, select qualified annotators, and perform quality inspection of annotation results. Our annotation document gives clear definitions and provides more than 10 examples for each judgment, to align the standard among annotators. Due to these efforts, the consistency of annotators in the binary results reaches 89.29% (images) and 89.33% (videos).
Annotation Analysis. Through specialized annotation, we obtain an image dataset containing 48k images and 3 million question-answer pairs, while a video dataset with 33k videos and 2 million pairs. More statistical analysis of the annotation results is in Appendix Section 7.2.
3.2 VisionReward Training
The complete training process of VisionReward and its application methodology during preference optimization are illustrated in  Fig. 3.
Fine-grained Visual Assessment. Specifically, we use CogVLM2 (hong2024cogvlm2) as the base model for image understanding, and CogVLM2-Video (hong2024cogvlm2) as the base model for video understanding. In terms of data, we have obtained millions of annotated binary question-answering pairs. Initially, we performed a balanced sampling on each binary question by addressing the imbalance between positive and negative examples, ensuring a roughly equal number of positive and negative instances associated with each binary question. Then we use balanced instruction tuning dataset consisting of binary questions to fine-tune base VLM.
Interpretable Preference Learning. After trained on fine-grained dataset, VisionReward can be adopt to judge any image or video by giving a series of binary response answers (“yes” or “no”) $\{A_{i}\}_{i=1}^{N}$, where $N$ represents the number of binary questions. We define reward of every binary question as $\{x_{i}\}_{i=1}^{N}$:
$x_{i}=\mathds{1}\left[A_{i}=\text{``yes''}\right].$ (3)
We construct a feature vector $X=(x_{1},\ldots,x_{N})$, and use a set of linear weights $W=(w_{1},\ldots,w_{N})$ to obtain the final reward $R$:
$R=\sum_{i=1}^{N}w_{i}\mathds{1}\left[A_{i}=\text{``yes''}\right].$ (4)
In order to learn linear weights $W$, we collect human preferences for pairs of $\{(X_{i},X_{j})\}$. Specifically, we compute the feature difference for each pair, given by $\Delta X=X_{i}-X_{j}$, and the corresponding label is assigned as $y=1$ or $y=0$ depending on the human preference. We then perform logistic regression $y=\Delta XW^{T}$ to learn linear weights $W$:
$\mathcal{L}(W)=$ $-\mathbb{E}\left[y\log\left(\sigma(\Delta XW^{T})\right)\right$.
$\left.+(1-y)\log\left(1-\sigma(\Delta XW^{T})\right)\right].$ (5)
By calculating dimension-specific scores through intra-dimensional weighting, VisionReward facilitates multi-dimensional preference prediction. We note dimensions as $\{\text{dim}_{k}\}_{k=1}^{K}$ where $\text{dim}_{k}$ contains questions belonging to the dimension. Then we define reward for certain dimension as:
$R({\text{dim}_{k}})=\sum_{i\in\text{dim}_{k}}w_{i}\mathds{1}\left[A_{i}=\text{``yes''}\right].$ (6)
3.3 Multi-Dimensional Preference Optimization
To empirically validate the model’s capacity, we leverage Direct Preference Optimization (DPO) (wallace2024diffusion) for Diffusion Models in our experiments, where VisionReward generates multi-dimensional preference pairs to guide the optimization process while maintaining inter-dimensional balance.
(a) Data analysis.
(b) DPO analysis.
Figure 4: (a) We sample 10,000 human preference pairs from Pick-a-Pic dataset and analyze score deviations across 18 sub-dimensions (represented by the average yes-proportion of checklist questions within each sub-dimension). (b) We show score deviations for images generated by SDXL after Diffusion-DPO, using the same 10,000 prompts.
Challenges. We replicate the Diffusion-DPO training procedure using SDXL (podell2023sdxl) on the Pick-a-Pic (kirstain2023pick) dataset, employing VisionReward for comprehensive data analysis and model evaluation. As demonstrated in  Fig. 4, both the preference data and optimized model exhibit biases across several fine-grained dimensions. These findings not only underscore VisionReward’s capability for fine-grained analysis but also emphasize the necessity for optimization approaches that account for multi-dimensional representation.
MPO: Insight and Solution. Compared to ordinary DPO method which select pairs using overall preference, we propose MPO-enhanced DPO that take account fine-grained multi-dimensional preference. For reward of two samples $R^{i}$ and $R^{j}$, we define $R^{i}$ as dominating $R^{j}$ if $R^{i}(\text{dim}_{k})\geq R^{j}(\text{dim}_{k})$ holds for every dimension $\text{dim}_{k}$. The key differences between the MPO strategy and standard DPO are:
  • Ordinary DPO: During DPO optimization, we directly select the pair based on the total reward $R$.
  • MPO-enhanced DPO: MPO strategy introduces an additional constraint: we only select pairs that $R^{i}$ dominates $R^{j}$, then proceed with standard DPO.
We analyze the effects of MPO in Section 4.3. The MPO strategy can also be applied to other algorithms, which we leave for future exploration.
Method Image Video
HPDv2 MonetBench GenAI-Bench MonetBench
tau diff∗∗ tau diff tau diff
task-specific discriminative models
ImageReward (xu2023imagereward) 74.0 48.8 56.5 48.4 72.1 55.8 58.4
PickScore (kirstain2023pick) 79.8 49.8 57.6 52.4 75.4 57.7 61.6
HPSv2 (wu2023human) 83.3 48.4 55.6 49.3 73.0 59.3 62.5
MPS (zhang2024learning) 83.5 44.2 50.7 46.9 67.6 55.8 58.9
generative models
GPT-4o (achiam2023gpt) 77.5 38.9 52.7 41.8 54.3 45.7 48.3
Gemini (team2024gemini) 60.7 27.4 55.1 46.9 61.7 52.2 56.8
VQAScore (lin2025evaluating) 69.7 49.4 56.5 45.2 68.0 56.1 59.5
VideoScore (he2024videoscore) 76.8 45.8 52.5 47.8 71.4 49.1 54.9
VisionReward (Ours) 81.7 51.8 59.5 51.8 74.4 64.0 72.1
Table 2: Preference accuracy on multiple dataset. Bold denotes the best score within the generative models, while underline signifies the best score among all categories. Tau means taking account of ties (deutsch2023ties), and diff∗∗ means dropping ties in labels (we drop ties both in labels and responses for GPT-4o and Gemini in diff∗∗ because too many ties are given by them).
4 Experiments
4.1 VisionReward for Text-to-Vision Evaluation
Dataset & Training Setting. After balanced sampling, we obtain 40,743 images and corresponding 97,680 judgment questions for training, leaving 6,910 images for subsequent validation and test. For videos, we obtain 28,605 videos and corresponding 89,473 judgment questions for training, with 3,080 videos reserved.
To fine-tune CogVLM2 (hong2024cogvlm2), we set a batch size of 64, a learning rate of 1e-6, and train for 1,500 steps. For CogVLM-Video, we set a batch size of 64, a learning rate of 4e-6, and train for 1,500 steps.
To learn linear weights for preference prediction, we sample human preference pairs and perform logistic regression. For images, We sample 44k pairs (24k from HPDv2 (wu2023human) and 20k from ImageRewardDB (xu2023imagereward)); and for videos, we sample prompts from VidProM (wang2024vidprom) and generate videos (using CogVideoX (yang2024cogvideox), VideoCrafter2 (chen2024videocrafter2) and OpenSora (opensora)), getting 1,795 annotated video pairs with preference.
To establish a comprehensive evaluation benchmark for both image and video generation, we construct MonetBench, which contains separate test sets for images and videos, each consisting of 1,000 prompts. More details are introduced in Appendix Section 13.
Figure 5: The accuracy of VisionReward improves as the number of binary questions increases. After masking weights from full regression, VisionReward maintains high performance.
Figure 6: Human evaluation results of DPO using different reward models or preference datasets. We require five annotators to comprehensively evaluate two samples and select the better one. VisionReward achieve the best performance.
Main Results: Preference Accuracy. Preference accuracy means the probability that a reward model has the same judgment as humans about which image is better. We use MonetBench to construct our test set for human preference, using SDXL (podell2023sdxl) to generate images and CogVideoX (yang2024cogvideox) / VideoCrafter2 (chen2024videocrafter2) / OpenSora (opensora) to generate videos, resulting in 500 pairs for image and 1,000 pairs for video. We employ annotators to assess the generated images using a preference rating scale from 1 to 5 (with 3 indicating no preference). The average preference score is used as the final preference label. We also take HPDv2 (wu2023human) and GenAI-Bench (jiang2024genai) as test set.
Table 2 shows that VisionReward obtains state-of-the-state results in multiple datasets. Notably, in video evaluation, image reward models demonstrate competitive performance when the video duration is within 2 seconds (GenAI-Bench). However, when the video duration reaches 6 seconds (MonetBench), only VisionReward is capable of accurately predicting human preference, being twice (22.1% over random) as high as the best (12.5% over random) among other methods. This indicates that dynamic information in longer videos poses a challenge for RMs, while VisionReward can effectively address this issue after fine-grained visual learning.
Ablation Study: Scalability of Question Scale.  Fig. 5 shows that accuracy of preference prediction exhibits significant improvement as the question scale increases, and masking minimal weights maintains performance (Details in Appendix  Section 8). Scalability validates the effects of decomposing preferences via fine-grained questions.
We do more analysis of VisionReward in Appendix  Section 8 and analysis of fine-grained results in  Section 11.
4.2 VisionReward for Preference Optimization
To evaluate the efficacy of VisionReward in preference optimization for visual generative systems, we conduct a series of comparative experiments against current state-of-the-art reward models and established preference datasets. We use SDXL as text-to-image base model and CogVideoX as text-to-video base model. The empirical results presented in  Fig. 6 demonstrate VisionReward’s superior performance, achieving statistically significant improvements in human preference metrics over competing approaches.
This section focuses on our experiments on preference optimization for text-to-video using different reward models. Details of experiments for text-to-image are provided in Appendix  Section 9.2 due to page limits.
Dataset & Training Settings. For our backbone model, we select CogVideoX-2B. The training prompts are sampled from VidProM (wang2024vidprom) (details in Appendix Section 9.1). To adapt these prompts for video generation, we have optimized them following guidelines from CogVideoX (yang2024cogvideox), which results in roughly 22,000 samples. We generate 4 videos for each prompt, and use VisionReward to score these videos and apply the MPO strategy to select approximately 9,400 effective preference pairs. In all our experiments, we maintain a batch size of 32, a learning rate of 5e-6, and employ 100 warmup steps followed by linear decay. We set the DPO parameter $\beta$ to 500. The MPO training process spans around 500 steps, equivalent to about 2 epochs. During training, we save a checkpoint every 40 steps and use a validation set split from the training set to pick the checkpoint with the highest reward.
Evaluation Settings. To comprehensively assess the MPO models, we have conducted both automatic and human evaluations. The automatic evaluation is conducted across various benchmarks, including VBench (huang2024vbench) and our Video-MonetBench. For VBench, we focus on commonly reported key metrics, including Human Action, Scene, Multiple Objects, and Appearance Style. In all these experiments, we utilize prompt optimization recommended in CogVideoX. Our baseline comparisons include the original CogVideoX-2B and DPO with VideoScore.
Methods    ActionHuman Scene    ObjectsMultiple    StyleAppear.
Original 98.20 55.60 68.43 24.20
VideoScore 97.60 56.25 68.66 23.96
VisionReward 98.40 57.57 71.54 24.02
Table 3: Evaluation results on VBench.
Methods Stability Dynamic Physics Preservation
Original 0.272 0.047 0.323 0.584
VideoScore 0.242 0.046 0.319 0.557
VisionReward 0.309 0.036 0.337 0.661
Table 4: Evaluation results on MonetBench.
Experimental Results. The main results are shown in  Table 3 and  Table 4. When compared to the original CogVideoX-2B, optimization with VisionReward significantly enhances model performance across these benchmarks. In contrast, optimization with VideoScore tends to degrade performance. The empirical evidence substantiates VisionReward’s advanced capacity for multi-dimensional optimization. (Case study in Appendix Section 9.3.)
4.3 Ablation Study of MPO
To comprehensively illustrate how MPO addresses the factor optimization bias inherent in DPO, We conduct experiments based on CogVideoX-5B. We set the threshold of the total score to 0.8 for DPO and 0.6 for MPO, ensuring that the number of pairs obtained through all three strategies is 5k. We use a batch size of 64, a learning rate of 2e-6, the DPO parameter $\beta$ of 500, and the training steps of 300. VisionReward is employed to evaluate scores across various dimensions, with the detailed results presented in  Table 5. Through the implementation of the MPO strategy, CogVideoX is optimized in such a manner that it avoids the degradation of certain factors (e.g., alignment), thereby achieving improved trade-offs, such as maintaining good preservation while avoiding excessively slow dynamic changes. The empirical evidence further substantiates VisionReward’s capacity for algorithm-agnostic preference alignment, as evidenced by comparative testing with other approaches like MaPO (hong2024margin).
Method Align. Quality Dynamic Physics Preserv. Overall
Original 1.733 0.660 0.053 0.344 0.653 4.303
DPO 1.697 0.680 0.034 0.356 0.741 4.515
DPO w/ MPO 1.766 0.688 0.042 0.356 0.721 4.573
MaPO 1.736 0.660 0.052 0.345 0.645 4.295
MaPO w/ MPO 1.737 0.656 0.055 0.349 0.649 4.321
Table 5: Ablation Study of MPO strategy. Scores are given by VisionReward on MonetBench.
For efficiency discussion, the experimental results in  Table 6 reveal the comparative effectiveness of the MPO and DPO methods. We analyze the pairs selected in perspective of $R^{i}$ dominating $R^{j}$” mentioned in  Section 3.3. These results suggest that MPO outperforms the DPO approach in terms of both efficiency and effectiveness. These findings highlight promising directions for future research in developing novel optimization algorithms and adapting VisionReward for multi-dimensional optimization.
Method #Dom. #Not-Dom. Reward
Original - - 4.303
DPO w/o MPO 3814 1456 4.515 (+0.212)
DPO w/ MPO 5028 0 4.573 (+0.270)
$\Delta$ (vs w/o MPO) +31.8% -100% +27.4%
Table 6: Comparison of MPO and DPO on MonetBench. “#Dom.” means the number of pairs that match the rule of $R^{i}$ dominating $R^{j}$”, while “#Not-Dom.” pairs not match.
5 Conclusion
We introduce VisionReward, a reward model for visual generation, which is fine-grained and multi-dimensional. By enabling Vision-Language Model (VLM) to perform binary assessments and applying linear summation with weighting coefficients derived from preference learning, VisionReward achieves highly accurate and interpretable. For visual generative optimization, VisionReward surpasses other reward models and enable multi-dimensional strategy.
6 Acknowledgments
The authors would like to thank Z.AI for sponsoring the computation resources used in this work.
Appendix
7 More Details of Annotation
7.1 Details of Annotation Design
To ensure the diversity of our annotation, we collect our annotated data from various sources, as presented in  Table 7. The prompt and example for data cleaning is shown in  Table 9.  Table 8 illustrates our preference dimensions and the checklist count. We identify 5 dimensions for text-to-image generation and expand to 9 dimensions for text-to-video.
Type Source #Samples #Checklist
Image ImageRewardDB (xu2023imagereward) 16K 1M
Pick-a-Pic (kirstain2023pick) 16K 1M
HPDv2 (wu2023human) 16K 1M
Video CogVideoX (yang2024cogvideox) 10K 0.6M
Open-Sora (opensora) 10K 0.6M
VideoCrafter2 (chen2024videocrafter2) 10K 0.6M
Panda-70M (chen2024panda) 3K 0.2M
Table 7: Statistics of source data and annotation.
Dimension #Sub-dimension #Checklist
Image Video Image Video
Alignment 1 1 1 4
Composition 5 1 13 2
Quality 5 4 14 14
Fidelity 5 3 25 9
Safety&Emotion 2 1 8 4
Stability - 5 - 12
Dynamic - 2 - 8
Physics - 1 - 4
Preservation - 2 - 7
Total 18 20 61 64
Table 8: Taxonomy of annotation for VisionReward.
(a) Text-to-image
(b) Text-to-video
Figure 7: Annotation statistics of different sub-dimensions.
SYSTEM Assume you are a model responsible for refining and polishing English expressions. You will receive an English prompt that may contain abbreviations or non-standard expressions. Your task is to standardize the expressions, and your output must be in pure English without any non-English characters. If the prompt is fragmented or difficult to understand, discard it by outputting “F”. Your output must strictly follow the format: each sentence should be on a single line, either as the rewritten prompt or a standalone “F”.
USER Here is the prompt you have received: [[PROMPT]]
INPUT Soft rays of light through the many different types of trees inside a forest, sunrise, misty, photorealistic, ground level, -neg "no large bodies of water" -ar 16:9 4K, -ar 16:9
OUTPUT The soft rays of light filter through the myriad types of trees within the forest at sunrise, creating a misty, photorealistic scene from ground level. Exclude any large bodies of water. The aspect ratio should be 16:9 in 4K resolution. Aspect ratio: 16:9.
Table 9: Prompt template and example for prompt cleaning.
7.2 Statistics of Annotation Result
Fig. 7 is the statistical results of the labeled data for images and videos. When compiling the statistics, higher labels indicate better performance for the image or video sub-dimension, while a label of 0 indicates neutrality. For video data, the original labels only had positive values and the neutral value was inconsistent. Hence, a neutral value was determined for each sub-dimension, and the original labels were adjusted by subtracting this neutral value to make 0 represent neutrality. In sub-dimensions such as Background, Face, and Hand, there might be cases where these elements are not present in the image or video. In such instances, ”Not Contain” is treated as a separate category for statistical purposes.
There are two main characteristics to note.
  • For most sub-dimensions, the distribution of options roughly follows a normal distribution, with the majority being ordinary, and the quantities of instances with extreme characteristics, either very good or very bad, are reduced. To assist the model in learning the features of each sub-dimension, we can impose a quantitative limit on the predominant options.
  • Certain sub-dimensions, such as the presence of hands, require a mask when predicting human preferences. This means that the sub-dimension should only be evaluated when the image indicates the presence of hands. We also annotate the sub-dimensions that require a mask and record the relevant counts.
8 More Details and Results of VisionReward
VisionReward comprises two steps: visual judgment and linear regression.
Visual Judgment Process. As the options for each sub-dimension are progressive, for any given option corresponding to a judgment question in the checklist, samples with an option greater than or equal to the given one are considered positive examples for the judgment question, whereas samples with an option less than the given one are considered negative. To balance the number of positive and negative examples for each binary question, we screen out any excess positive and negative examples for each question, ensuring that the number of positive and negative examples used for training is balanced. For alignment, we use different methods on images and videos. For images, we use VQAScore (lin2025evaluating) as an alignment judgment. For videos, we train five levels of judgment for VisionReward.
Main Results: Accuracy on Judgment. To evaluate the effectiveness of judgment learning, we construct a visual quality QA set to assess the visual assessment capabilities of VisionReward compared to other VLMs. We collect 1,364 test cases for images involving 14 types of questions across 4 dimensions, and additionally 1,308 cases covering 8 types of questions across 4 dimensions related to dynamic content in videos. To ensure the generality of these questions, we combine adjacent degrees in the checklists under each sub-dimension, enhancing distinctiveness and minimizing incidental subjectivity.  Table 11 demonstrates VisionReward’s superiority in judging visual quality over existing generalized multimodal LLMs, which remains a challenge for generalized multimodal LLMs.
Composition Quality Fidelity Safety
97.9 98.2 98.3 99.1
Stability Dynamic Physics Preservation
97.4 99.9 88.2 99.8
Table 10: Consistency of VisionReward in each dimension.
Main Results: Consistency on Judgment. As questions corresponding to each sub-dimension assess varying degrees of a particular factor, it’s important to measure consistency of VisionReward across multiple questions of the same sub-dimension. Consistency measures the likelihood that the model provides consistent responses across a series of judgments concerning this factor.  Table 10 shows that VisionReward has high consistency (more than 97 %) in most (7 of 8) dimensions.
Algorithm 1 Iterative Regression with Weight Masking
0: Dataset of human preferences $\mathcal{D}=\{(\mathbf{X}_{i},\mathbf{X}_{j},y)\}$, where each pair $(\mathbf{X}_{i},\mathbf{X}_{j})$ represents feature vectors of binary responses and $y\in\{0,1\}$ is the human preference label
1:Initialization: Initialize linear weights $\mathbf{w}=[w_{1},\ldots,w_{n}]$
2: Initialize convergence criterion $\text{diff}\leftarrow\infty$
3:while $\text{diff}>\epsilon$ do
4:  $\textbf{w}_{\text{old}}\leftarrow\textbf{w}$
5:  for each $(\mathbf{X}_{i},\mathbf{X}_{j},y)$ in $\mathcal{D}$ do
6:   $\Delta\mathbf{X}\leftarrow\mathbf{X}_{i}-\mathbf{X}_{j}$
7:   $\hat{y}\leftarrow\sigma(\Delta\mathbf{X}^{T}\mathbf{w})$
8:   $\nabla_{\textbf{w}}\text{Loss}=(\hat{y}-y)\Delta\mathbf{X}$
9:   $\mathbf{w}\leftarrow\mathbf{w}-\alpha\nabla_{\textbf{w}}\text{Loss}$
10:  end for
11:  Mask negative weights $\mathbf{w}\leftarrow\mathbf{w}\odot(\mathbf{w}>0)$
[…]
13:end while
13: Trained weights $\mathbf{w}$
Method Image Video
Composition Quality Fidelity Safety Stability Dynamic Physics Preservation
LLaVa 59.9 65.7 59.8 64.4 52.5 53.8 50.6 47.5
CogVLM2 (hong2024cogvlm2) 65.8 67.1 53.1 74.7 49.3 57.1 51.2 47.8
GPT-4o (achiam2023gpt) 73.1 62.7 61.9 70.1 57.9 69.1 62.4 58.8
Gemini (team2024gemini) 69.4 59.9 59.7 74.9 58.1 71.1 58.1 59.6
VisionReward (Ours) 78.8 81.1 80.9 83.9 64.8 75.4 68.1 72.0
Table 11: Accuracy of VisionReward and other vision-language models (VLMs) on vision quality questions constructed from our annotation. We test LLaVA-v1.5-7B (liu2023llava) for image and LLava-Next-Video-34B (li2024llava) for video.
Weight Masking. In the linear regression step, we learn the correlation between human preferences and the results of visual judgment. In our design, if the result of the visual judgment is “yes”, human preference improves. We examine the correlation between human preference and each judgment result, and the numerical results indicate a positive correlation. However, in linear regression, we observe that some coefficients corresponding to the judgment results were negative. This is because there are correlations among the judgment results themselves. To enhance the robustness of the regression outcomes, we employ an iterative masking algorithm during the regression.
Ablation Study: Impact of Training Set Size. We conduct experiments with varying sizes of the training set to investigate its influence on regression performance. As demonstrated in  Table 12, the accuracy improves monotonically as the training set size increases up to 4,000 samples, beyond which the accuracy stabilizes.
Size 200 500 1k 2k 4k 8k
Acc. 77.6 80.3 80.6 80.9 81.3 81.3
Table 12: Accuracy on HPDv2-test for different sizes of train set.
9 More Details and Results of MPO
9.1 Details of Prompt for MPO
We employ the prompt filtering approach proposed by UniFL (zhang2024uniflimprovestablediffusion) to curate our dataset. This strategy comprises two pivotal steps:
  • Semantic-Based Filtering: Utilizing an existing scene graph parser (8954449), we evaluate the semantic richness of prompts by analyzing the number of subjective and objective relationships. Prompts with fewer than one meaningful relationship are filtered out to reduce noise data.
  • Cosine Similarity-Based Selection: Following the semantic filtering process, we apply a cosine similarity-based iterative selection mechanism. By maintaining a maximum similarity threshold of 0.8 between any two prompts, we ensure dataset diversity and effectively eliminate redundant entries.
9.2 VisionReward for Text-to-Image Optimization
Dataset & Training Settings. We strategically sample 63,165 prompts from existing datasets and generate 8 images per prompt using SDXL (this procedure theoretically produces 1.76M text-image pairs). Employing the MPO algorithm, we obtain 760k dominant pairs with 63,069 unique prompts. For comparison, we use HPSv2 with a threshold of 0.0015, getting 770k pairs with 63,107 unique prompts from the same source. We also compare with human annotated pairs, sampling 780k human preference pairs with 57,674 unique prompts from Pick-a-Pic v2 dataset.
We maintain consistent training parameters and dataset sizes across all experiments to ensure fair comparison. For all three experiments, we used an effective batch size of 256 (with GAS set to 4 and train batch size set to 1), set $\beta$ to 5000, and a learning rate of 5e-9 (before scaling). We employ a constant warmup strategy with 100 steps and the training is conducted over 3,000 steps (approximately 1 epoch).
Evaluation Settings. We conducted both automatic and human evaluation on DrawBench (saharia2022photorealistic). Automatic evaluation includes multiple metrics such as human preference RMs, CLIP (radford2021learning) and LAION-Aesthetic (schuhmann2022laion).
Experimental Results. Main results are demonstrated in  Table 13 and  Table 14. VisionReward gets leading results across multiple machine metrics and achieves significant improvements across all four dimensions of VisionReward.
Methods CLIP Aes HPSv2 PickScore
Original 0.273 5.463 0.282 22.25
Pick-a-Pic 0.279 5.511 0.286 22.45
HPSv2 0.277 5.599 0.292 22.58
VisionReward 0.279 5.612 0.289 22.61
Table 13: Evaluation results of multiple metrics on DrawBench.
Methods Composition Quality Fidelity Safety
Original 0.755 0.550 0.009 -0.008
Pick-a-Pic 0.765 0.588 0.009 -0.009
HPSv2 0.874 0.630 0.010 -0.004
VisionReward 0.894 0.670 0.017 -0.001
Table 14: Evaluation results on MonetBench.
9.3 More Results of MPO
Case Study.  Fig. 8 shows MPO cases for text-to-image, while  Fig. 9 and  Fig. 10 show MPO cases for text-to-video. MPO fine-tuned model surpasses the original model in multiple aspects and also outperforms other scoring methods.
Training Curve. Fig. 11 shows the variation of the dimensional scores during the MPO process with respect to the number of training samples. The results demonstrate that the MPO method enables the model to avoid trade-offs during training, thereby achieving simultaneous improvements across various sub-dimensions. In contrast, the DPO (wallace2024diffusion) method fails to achieve this level of concurrent enhancement.
Ablation Study: Different Strategy for MPO. In using the MPO strategy, we define the way in which $R^{i}$ dominates $R^{j}$ (given two images $x^{i}$ and $x^{j}$) to select pairs. The definition of “dominate” includes at least three methods for different reward objective: score of each dimension, score of each sub-dimension, and score of each binary question. To investigate the impact of different definitions of “dominate”, we conduct experiments based on CogVideoX-5B. Specifically, we employ three different strategies, setting the threshold of the total score to 0.6, 0.5, and 0.4 respectively, ensuring that the number of pairs obtained through all three strategies is 5k. We use a batch size of 64, a learning rate of 2e-6, the DPO parameter $\beta$ of 500, and the training steps of 300. After training, we compare the evaluation results of VisionReward on MonetBench.  Table 15 shows that using the score for each dimension as objective yields the best results.
Methods Baseline Dimension Sub-dimension Question
Reward 4.303 4.573 4.539 4.514
Table 15: Score of VisionReward after different strategies of MPO. Dimension: “dominate” based on score of each dimension. Sub-dimension: “dominate” based on score of each sub-dimension. Question: “dominate” based on score of each binary question.
Figure 8: Qualitative result of MPO in text-to-image.
Figure 9: Qualitative result of MPO in text-to-video.
Figure 10: Qualitative result of MPO in text-to-video.
(a) Overall Score
(b) Composition Score
(c) Fidelity Score
(d) Alignment Score
(e) Quality Score
(f) Safety & Emotion Score
Figure 11: Variation of dimensional scores during the MPO process with respect to the number of training samples.
10 Details of Fine-Grained Questions
Table 16 and  Table 17 show the annotation taxonomy of images, while  Table 18 and  Table 19 show for videos.
11 More Results of Fine-Grained Design
Weight and Accuracy of Checklist. We curate separate test sets of images and videos from outside the training set to evaluate the accuracy of judgment questions. The test set comprises 1,209 images and 1,000 videos, respectively. We report the accuracy of judgment questions (Cf.  Table 20 and  Table 21 for text-to-image,  Table 22 and  Table 23 for text-to-video). As a reference, we specifically record the linear weights obtained from linear regression on human preference data as well as the Spearman rank correlation coefficient between human preference and the results of each judgment question.
Correlation of Sub Dimensions. To mine the correlation between the sub-dimensions after preference decoupling, we show the correlation coefficients between the sub-dimensions in a heat map (Cf.  Fig. 12).
Dimension Sub-dimension Option Checklist
Composition Symmetry symmetrical Is the image symmetrical?
ordinary Does the image avoid asymmetry?
asymmetrical
Composition Object pairing coordinated Are the objects well-coordinated?
ordinary Does the image avoid poorly coordinated objects?
uncoordinated
Composition Main object prominent Is the main subject prominent?
ordinary Does the image avoid an unclear main subject?
prominent
Composition Richness very rich Is the image very rich?
rich Is the image rich?
ordinary Is the image not monotonous?
monotonous Is the image not empty?
empty
Composition Background beautiful Is the background beautiful?
somewhat beautiful Is the background somewhat beautiful?
ordinary Is there a background?
no background
Quality Clarity very clear Is the image very clear?
clear Is the image clear?
ordinary Does the image avoid being blurry?
blurry Does the image avoid being completely blurry?
completely blurry
Quality Color Brightness bright Are the colors bright?
ordinary Are the colors not dark?
dark
Quality Color Aesthetic beautiful colors Are the colors beautiful?
ordinary colors Are the colors not ugly?
ugly colors
Quality Lighting Distinction very distinct Is the lighting and shadow very distinct?
distinct Is the lighting and shadow distinct?
ordinary Is there lighting and shadow?
no lighting
Quality Lighting Aesthetic very beautiful Are the lighting and shadows very beautiful?
beautiful Are the lighting and shadows beautiful?
ordinary Is there lighting and shadow?
no lighting
Table 16: Annotation taxonomy and checklist details for text-to-image evaluation. (part 1)
Dimension Sub-dimension Option Checklist
Fidelity Detail reality realistic Are the image details realistic?
neutral Do the image details avoid being unrealistic?
unrealistic Do the image details avoid being very unrealistic?
very unrealistic Do the image details avoid being greatly unrealistic?
greatly unrealistic
Fidelity Detail refinement very refined Are the image details very exquisite?
refined Are the image details exquisite?
ordinary Do the image details avoid being coarse?
rough Do the image details avoid being very coarse?
very rough Does the image avoid being hard to recognize?
indistinguishable Does the image avoid being fragmented?
fragmented
Fidelity Body no errors Is the human body in the image completely correct?
neutral Does the human body in the image avoid errors?
some errors Does the human body in the image avoid obvious errors?
obvious errors Does the human body in the image avoid serious errors?
serious errors Is there a human body in the image?
no human figure
Fidelity Face very beautiful Is the human face very beautiful?
beautiful Is the human face beautiful?
normal Does the human face avoid errors?
some errors Does the human face avoid serious errors?
serious errors Is there a human face in the image?
no human face
Fidelity Hands perfect Are the human hands perfect?
mostly correct Are the human hands essentially correct?
minor errors Do the human hands avoid obvious errors?
obvious errors Do the human hands avoid serious errors?
serious errors Are there human hands in the image?
no human hands
Safety & Emotion Emotion very positive Can the image evoke a very positive emotional response?
positive Can the image evoke a positive emotional response?
ordinary Does the image avoid evoking a negative emotional response?
negative Does the image avoid evoking a very negative emotional response?
very negative
Safety & Emotion Safety safe Is the image completely safe?
neutral Is the image harmless?
potentially harmful Does the image avoid obvious harmfulness?
harmful Does the image avoid serious harmfulness?
very harmful
Table 17: Annotation taxonomy and checklist details for text-to-image evaluation. (part 2)
Dimension Sub-dimension Option Checklist
Alignment Alignment meet 100% Does the video meet all the requirements stated in the text ”[[prompt]]”?
meet 80%-100% Does the video meet most of the requirements stated in the text ”[[prompt]]”?
meet 60%-80% Does the video meet some of the requirements stated in the text ”[[prompt]]”?
meet 40%-60% Does the video not completely fail to meet the requirements stated in the text ”[[prompt]]”?
meet 0-40%
Composition Composition good Is the composition aesthetically pleasing?
normal Does the composition have no obvious flaws?
bad
Quality Color very beautiful Are the colors exceptionally beautiful?
beautiful Are the colors beautiful?
normal Are the colors not unattractive?
unattractive Are the colors not significantly unattractive?
very unattractive
Quality Lighting Accurate good Is the lighting perfectly accurate?
normal Does the lighting have no obvious errors?
bad Is there any lighting present?
no lighting
Quality Lighting Aes very good Is the lighting exceptionally beautiful?
good Is the lighting beautiful?
normal Is the lighting not unattractive?
bad
Quality Clear very clear Is it very clear?
clear Is it clear?
normal Is it not blurry?
blurry Is it not completely blurry?
completely blurry
Fidelity Detail Refinement very refined Are the details very refined?
refined Are the details refined?
normal Are the details not rough?
rough Are the details not significantly rough?
very rough
Fidelity Movement Reality good Is the object’s movement completely realistic?
normal Does the object’s movement have no obvious realism issues?
bad
Fidelity Letters good Are all the letters correct?
normal Do the letters have no obvious errors?
bad Are there any letters present?
no letter
Safety Safety 100% safe Is the video content safe?
80%-100% safe Is the video content definitely free of harmful material?
60%-80% safe Does the video content contain no harmful material?
40%-60% safe Does the video content contain no extremely harmful material?
0-40% safe
Table 18: Annotation taxonomy and checklist details for text-to-video evaluation. (part 1)
Dimension Sub-dimension Option Checklist
Stability Movement smoothness good Is the smoothness of the object’s movement good?
normal Does the smoothness of the object’s movement have no obvious issues?
bad
Stability Image quality stability very stable Is the image quality very stable?
stable Is the image quality stable?
normal Is the image quality not unstable?
unstable Is the image quality free of noticeable instability?
very unstable
Stability Focus good Is the focus aesthetically pleasing?
normal Does the focus have no obvious flaws?
bad
Stability Camera movement good Is the camera movement aesthetically pleasing?
normal Does the camera movement have no obvious flaws?
bad
Stability Camera stability stable Is the camera stable?
normal Is the camera not unstable?
unstable
Preservation Shape at beginning completely accurate Is the shape of the object at the beginning of the video completely accurate?
no errors Does the shape of the object at the beginning have no obvious errors?
not chaotic Is the shape of the object at the beginning not chaotic?
flawed
Preservation Shape throughout perfectly maintained Is the shape of the object perfectly maintained throughout the video?
no issues Does the shape of the object have no obvious issues throughout the video?
normal Does the shape of the object generally have no major issues throughout the video?
not chaotic Is the shape of the object not chaotic throughout the video?
flawed
Dynamic Object Motion dynamic highly dynamic Is the object’s motion highly dynamic?
dynamic Is the object’s motion dynamic?
normal Is the object’s motion not minimal?
not static Is the object’s motion not static?
static
Dynamic Camera motion dynamic highly dynamic Is the camera motion highly dynamic?
dynamic Is the camera motion dynamic?
not minimal Is the camera motion not minimal?
not static Is the camera motion not static?
static
Physics Physics law full compliance Does it fully comply with the laws of physics?
partial compliance Does it partially comply with the laws of physics?
no obvious violations Does it have no obvious violations of the laws of physics?
physical world Is the video content part of the physical world?
non-compliance
Table 19: Annotation taxonomy and checklist details for text-to-video evaluation. (part 2)
ID Checklist Acc $\rho$ Weight
1 Is there a human body in the image? 93.13 0.090 mask
2 Is there a human face in the image? 96.20 0.110 mask
3 Are there human hands in the image? 93.30 0.022 mask
4 Is the image symmetrical? 79.98 0.104 0.069
5 Does the image avoid asymmetry? 71.30 0.236 0.102
6 Are the objects well-coordinated? 58.31 0.138 0.000
7 Does the image avoid poorly coordinated objects? 68.24 0.204 0.000
8 Is the main subject prominent? 86.27 0.210 0.131
9 Does the image avoid an unclear main subject? 77.75 0.258 0.070
10 Is the image very rich? 80.40 0.084 0.056
11 Is the image rich? 65.84 0.138 0.044
12 Is the image not monotonous? 77.01 0.271 0.211
13 Is the image not empty? 99.67 0.205 0.583
14 Is the background beautiful? 72.70 -0.019 0.000
15 Is the background somewhat beautiful? 67.26 0.021 0.000
16 Is there a background? 84.86 0.079 mask
17 Is the image very clear? 63.85 0.111 0.051
18 Is the image clear? 62.03 0.170 0.068
19 Does the image avoid being blurry? 88.92 0.284 0.065
20 Does the image avoid being completely blurry? 97.11 0.282 0.032
21 Are the colors bright? 63.69 0.098 0.076
22 Are the colors not dark? 82.88 0.141 0.077
23 Are the colors beautiful? 65.84 0.115 0.000
24 Are the colors not ugly? 74.77 0.232 0.042
25 Is the lighting and shadow very distinct? 75.45 -0.043 0.000
26 Is the lighting and shadow distinct? 58.37 0.035 0.000
27 Is there lighting and shadow? 75.93 0.108 mask
28 Are the lighting and shadows very beautiful? 80.47 -0.055 0.000
29 Are the lighting and shadows beautiful? 71.99 -0.026 0.000
30 Can the image evoke a very positive emotional response? 82.63 0.068 0.051
31 Can the image evoke a positive emotional response? 63.94 0.117 0.000
32 Does the image avoid evoking a negative emotional response? 76.01 0.179 0.000
33 Does the image avoid evoking a very negative emotional response? 91.56 0.117 0.000
34 Are the image details very exquisite? 74.03 0.078 0.010
35 Are the image details exquisite? 71.79 0.091 0.000
36 Do the image details avoid being coarse? 68.73 0.215 0.000
37 Do the image details avoid being very coarse? 84.62 0.247 0.000
38 Does the image avoid being hard to recognize? 87.34 0.267 0.017
39 Does the image avoid being fragmented? 85.36 0.288 0.115
40 Are the image details realistic? 63.85 0.099 0.000
Table 20: Accuracy, spearman correlation, and linear weights of VisionReward in text-to-image. (Part 1)
ID Checklist Acc $\rho$ Weight
41 Do the image details avoid being unrealistic? 63.94 0.140 0.000
42 Do the image details avoid being very unrealistic? 74.19 0.156 0.000
43 Do the image details avoid being greatly unrealistic? 83.62 0.177 0.000
44 Is the human body in the image completely correct? 61.31 0.063 0.082
45 Does the human body in the image avoid errors? 59.02 0.129 0.000
46 Does the human body in the image avoid obvious errors? 82.57 0.135 0.055
47 Does the human body in the image avoid serious errors? 90.83 0.121 0.030
48 Is the human face very beautiful? 65.50 -0.046 0.000
49 Is the human face beautiful? 56.88 -0.006 0.000
50 Does the human face avoid errors? 57.61 0.113 0.031
51 Does the human face avoid serious errors? 91.56 0.132 0.077
52 Are the human hands perfect? 90.18 -0.015 0.072
53 Are the human hands essentially correct? 25.84 0.059 0.000
54 Do the human hands avoid obvious errors? 37.98 0.066 0.000
55 Do the human hands avoid serious errors? 77.26 0.048 0.000
56 Is the image completely safe? 78.74 0.118 0.000
57 Is the image harmless? 86.44 0.106 0.000
58 Does the image avoid obvious harmfulness? 92.39 0.109 0.012
59 Does the image avoid serious harmfulness? 92.80 0.092 0.015
60 Does the image show ”[[prompt]]”? - 0.297 2.354
Table 21: Accuracy, spearman correlation, and linear weights of VisionReward in text-to-image. (Part 2)
ID Checklist Acc $\rho$ Weight
1 Does the video meet all the requirements stated in the text ”[[prompt]]”? 69.5 0.315 0.954
2 Does the video meet most of the requirements stated in the text ”[[prompt]]”? 72.9 0.303 0.252
3 Does the video meet some of the requirements stated in the text ”[[prompt]]”? 72.9 0.281 0.000
4 Does the video not completely fail to meet the requirements stated in the text ”[[prompt]]”? 78.7 0.320 1.142
5 Is the composition aesthetically pleasing? 50.8 0.263 0.035
6 Does the composition have no obvious flaws? 90.4 0.239 0.025
7 Is the focus aesthetically pleasing? 49.8 0.232 0.000
8 Does the focus have no obvious flaws? 91.6 0.246 0.000
9 Is the camera movement aesthetically pleasing? 76.2 0.012 0.000
10 Does the camera movement have no obvious flaws? 97.3 0.142 0.126
11 Are the colors exceptionally beautiful? 46.5 0.214 0.000
12 Are the colors beautiful? 50.1 0.217 0.000
13 Are the colors not unattractive? 82.2 0.225 0.000
14 Are the colors not significantly unattractive? 88.6 0.202 0.032
15 Is the lighting perfectly accurate? 51.9 0.346 0.163
16 Does the lighting have no obvious errors? 86.2 0.259 0.217
17 Is there any lighting present? 87.8 0.215 0.020
18 Is the lighting exceptionally beautiful? 65.1 0.212 0.136
19 Is the lighting beautiful? 55.8 0.240 0.096
20 Is the lighting not unattractive? 83.5 0.280 0.155
Table 22: Accuracy, spearman correlation, and linear weights of VisionReward in text-to-video. (Part 1)
ID Checklist Acc $\rho$ Weight
21 Is the shape of the object at the beginning of the video completely accurate? 63.0 0.292 0.129
22 Does the shape of the object at the beginning have no obvious errors? 76.3 0.274 0.099
23 Is the shape of the object at the beginning not chaotic? 91.3 0.256 0.188
24 Is the shape of the object perfectly maintained throughout the video? 54.2 0.300 0.184
25 Does the shape of the object have no obvious issues throughout the video? 68.8 0.267 0.000
26 Does the shape of the object generally have no major issues throughout the video? 84.5 0.259 0.000
27 Is the shape of the object not chaotic throughout the video? 93.5 0.240 0.264
28 Is the object’s motion highly dynamic? 78.0 -0.079 0.000
29 Is the object’s motion dynamic? 69.0 -0.024 0.000
30 Is the object’s motion not minimal? 71.2 -0.009 0.000
31 Is the object’s motion not static? 66.5 -0.014 0.000
32 Is the camera motion highly dynamic? 86.9 -0.054 0.112
33 Is the camera motion dynamic? 80.6 -0.062 0.000
34 Is the camera motion not minimal? 72.1 -0.061 0.052
35 Is the camera motion not static? 58.1 -0.059 0.000
36 Is the smoothness of the object’s movement very good? 59.8 0.263 0.026
37 Does the smoothness of the object’s movement have no obvious issues? 61.6 0.139 0.000
38 Is the object’s movement completely realistic? 66.8 0.338 0.439
39 Does the object’s movement have no obvious realism issues? 69.2 0.235 0.000
40 Is it very clear? 52.1 0.261 0.000
41 Is it clear? 51.0 0.290 0.000
42 Is it not blurry? 81.8 0.271 0.000
43 Is it not completely blurry? 93.1 0.226 0.000
44 Is the image quality very stable? 43.1 0.313 0.269
45 Is the image quality stable? 61.2 0.294 0.000
46 Is the image quality not unstable? 79.0 0.277 0.000
47 Is the image quality free of noticeable instability? 87.6 0.247 0.000
48 Is the camera very stable? 54.2 0.197 0.000
49 Is the camera not unstable? 83.5 0.267 0.000
50 Are the details very refined? 73.0 0.324 0.429
51 Are the details relatively refined? 62.3 0.331 0.000
52 Are the details not rough? 74.2 0.302 0.008
53 Are the details not significantly rough? 89.2 0.271 0.128
54 Are all the letters correct? 87.3 0.114 0.058
55 Do the letters have no obvious errors? 86.8 0.115 0.000
56 Are there any letters present? 89.7 0.104 0.145
57 Does it fully comply with the laws of physics? 36.6 0.254 0.000
58 Does it partially comply with the laws of physics? 66.7 0.248 0.000
59 Does it have no obvious violations of the laws of physics? 77.4 0.231 0.000
60 Is the video content part of the physical world? 86.6 0.231 0.394
61 Is the video content safe? 92.8 0.000 0.000
62 Is the video content definitely free of harmful material? 94.3 0.000 0.000
63 Does the video content contain no harmful material? 97.7 0.000 0.000
64 Does the video content contain no extremely harmful material? 100.0 0.000 0.000
Table 23: Accuracy, spearman correlation, and linear weights of VisionReward in text-to-video. (Part 2)
Figure 12: Correlation heatmap of text-to-video sub dimensions.
12 Limitations
This section outlines several limitations of the VisionReward framework that emerged during its development.
Supported Video Frame Rate and Length. While training datasets of VisionReward contain videos up to 6 seconds in duration at 4 frames per second (fps), this may be insufficient for evaluating the outputs of next-generation video generation models, which are capable of producing longer and more complex content. Therefore, extending the model’s capacity to handle longer video sequences is a critical direction for future work.
Leverage of Foundation Model. The model’s performance is inherently tied to the capabilities of its base Vision-Language Model (VLM). Our current approach utilizes a question-answering (QA) mechanism to tap into the VLM’s foundational knowledge. However, with the rise of sophisticated reasoning models, a more effective avenue for enhancing our reward model would be to integrate explicit reasoning abilities. This represents a key area for future investigation to move beyond foundational understanding towards more complex evaluation.
13 More Details of MonetBench
Type Image Video
Content People, Objects, Animals, Story, Human Activity,
Architecture, Landscape, Artificial Scene, Others,
Vehicles, Plants, Food, Natural Animal Activity,
Others, Scenes Physical Phenomena
Challenge Unreal, Style, History, Material, Angle and Lens,
Fine-grained Detail, Emotional Expression,
Color, Famous Character, Color/Tone, Surreal,
Normal, Famous Places, World Knowledge,
Writing, Complex Combo, Special Effects, Text,
Positional, Counting, Spatial Relationship,
Camera Movement,
Logical Consistency,
Style, Temporal Speed
Table 24: Content and Challenge of MonetBench.
Image-MonetBench Construction. We first establish our dataset foundation by collecting 4,038 seed prompts through strategic sampling from established datasets (1,000 from ImageRewardDB (xu2023imagereward), 1,000 from HPDv2 (xu2023imagereward), and 2,038 from Pick-a-Pic (kirstain2023pick)). Through systematic analysis of these prompts, we identify nine fundamental visual elements as content categories and twelve distinct aspects of generation complexity as challenge categories, maintaining the categorical distributions observed in the source datasets.
Video-MonetBench Construction. For video prompt evaluation, we initially sample 20,000 prompts from the VproM (wang2024vidprom) dataset, which are filtered to 13,342 valid entries after removing duplicates and invalid content. Our video classification system comprises seven content categories reflecting different video scenarios, and thirteen challenge categories capturing various technical and creative aspects of video generation.
Prompt Generation and Filtering. To ensure benchmark quality and diversity, we employ ChatGLM (glm2024chatglmfamilylargelanguage) to generate 1,000 new prompts for each benchmark following the established category distributions. Each generated prompt undergoes a three-stage filtering process: (1) Rouge-L (lin-2004-rouge) similarity checking for textual diversity, (2) semantic filtering with a cosine similarity threshold of 0.9, and (3) proportional sampling to maintain the intended category distributions.
Table 24 shows content and challenge categories for MonetBench. The resulting benchmark achieves balanced coverage across all categories while maintaining high standards of prompt diversity and quality. This carefully crafted multi-dimensional design enables comprehensive evaluation of visual reward models across both fundamental content types and various generation challenges.
For experimental efficiency, we provide a condensed version by randomly sampling 500 prompts from each benchmark while preserving the categorical distribution.
Details of Classification Proportions. After completing the design of the comprehensive two-dimensional classification framework, we utilized ChatGLM to categorize each prompt in the dataset across content and challenge dimensions. We then calculated the proportions of different classification labels for content and challenges. The content and challenge categories and their respective examples are summarized in Tables 25 and 26. Based on these proportions, we used ChatGLM to construct Benchmark prompts (all prompts were generated by ChatGLM, not directly sampled from the dataset). During the construction process, we specified the investigation direction and randomly sampled four ”seed prompts” from the categorized prompts to generate new, higher-quality prompts with ChatGLM. This synthesis approach produced two benchmark datasets, containing 1,000 and 1,007 meticulously crafted prompts, respectively, preserving the statistical characteristics of the original data.
The final datasets provide balanced and comprehensive coverage of content and challenge categories. Table 27 lists the specific content and challenge categories with detailed descriptions and example prompts, providing a clear understanding of the dataset’s composition. The structured methodology ensures the datasets’ diversity and alignment with real-world visual generation requirements, enabling nuanced benchmarking of visual models.
Image Video
Type Ratio Count Type Ratio Count
People 8 286 Story 5 265
Objects 4 143 Human Activity 4 212
Animals 4 143 Artificial Scene 3 159
Architecture 4 143 Natural Scenes 3 159
Others 2 72 Animal Activity 2 106
Landscape 2 72 Physical Phenomena 1 53
Vehicles 2 71 Other 1 53
Plants 1 35
Food 1 35
Table 25: Content Categories for Image and Video
Image Video
Type Ratio Count Type Ratio Count
Unreal 8 187 Style 13 465
Style & Format 8 187 Material/Texture 8 292
Fine-grained Detail 8 186 Emotional Expr. 7 249
Color 4 93 Color/Tone 7 261
Famous Character 4 93 World Knowledge 5 192
History & Culture 4 93 Special Effects 5 183
Normal 2 46 World Knowledge 4 192
Writing 1 23 Spatial Relat. 4 136
Complex Combo 1 23 Camera Move. 4 153
Famous Places 1 23 Surreal 3 108
Positional 1 23 Logical Consist. 2 116
Counting 1 23 Temporal Speed 1 66
Text 1 46
Table 26: Challenge Categories for Image and Video
Categorie Description Example Prompt
Content
Human Activity Descriptions about daily human activities, sports, performing arts, and professional skills. A family enjoying a picnic in a park, children playing soccer.
Animal Activity Descriptions about wild animals, domestic pets, and interactions between animals. A group of dolphins jumping out of the water.
Natural Scenes Descriptions about weather changes, geological events, and astronomical phenomena. A thunderstorm with lightning striking the ground.
Artificial Scenes Descriptions about cityscapes, interiors of buildings, vehicles, and industrial production. A bustling city street with traffic and pedestrians.
Physical Phenomena Descriptions about physical occurrences like candle burning, ice melting, glass breaking, and explosions. A glass shattering in slow motion.
Story Descriptions about coherent narratives based on a story or fantasy rather than a single scene or activity. Alice, a young girl, falls down a rabbit hole into a wonderland full of fantastical creatures and adventures.
Other Descriptions about various contents that do not fit into the other specified categories. Various clips of miscellaneous activities not fitting into other categories.
Challenge
Style Descriptions about artistic styles such as realistic, cyberpunk, and animated. A futuristic city with neon lights and flying cars, portrayed in a cyberpunk style.
Color/Tone Descriptions about color schemes like warm tones, cool tones, monochrome, and high saturation. A serene landscape in warm, golden tones during sunset.
Camera Movement Descriptions about different camera movements, including fixed, panning, zooming, tracking, and aerial shots. A drone shot capturing a bird’s eye view of a mountain range.
Special Effects Descriptions about special effects such as particle effects, lighting effects, and transitions. Fireworks exploding with sparkling particle effects.
Material/Texture Descriptions about materials and textures like metal, wood, glass, and fabric. Close-up shot of rain droplets on a glass window.
Surreal Descriptions about dreamlike, fantastical, or non-realistic elements. A dreamlike scene with floating islands in the sky.
Temporal Speed Descriptions about different speeds, including slow motion, normal speed, fast motion, and time reversal. Slow-motion capture of a hummingbird in flight.
Spatial Relationships Descriptions about the spatial arrangement of objects, their sizes, occlusions, and perspectives. A house of cards being built, showing each layer’s spatial arrangement.
World Knowledge Descriptions about physical laws, famous landmarks, historical events, and renowned personalities. A documentary about the pyramids of Egypt.
Logical Consistency Descriptions about ensuring logical relationships among events, timelines, and spatial layouts. A mystery story where clues are pieced together logically.
Emotional Expression Descriptions about expressions of emotions such as joy, sorrow, fear, and surprise. A close-up of a person expressing joy after receiving good news.
Text Descriptions about incorporating textual elements dynamically within the footage. An animated title sequence with dynamic text effects.
Table 27: Video classification standards with example prompts.